### v3

@@ III-A Verification of Train Schedules on ETCS L2 HTD Layouts
Assume that the given monotone 3-SAT-(2,2$\leq$ 2$\leq$ 2≤ 2 , ≤ 2) instance is feasible. From any particular solution, define $A:=\{j\colon x_{j}\text{ is true}\}$ and $A^{C}:=\{1,\ldots,m\}-A=\{j\colon x_{j}\text{ is false}\}$. For $1\leq i\leq k$, define $\tilde{x}(i):=\min\{j\colon j\in A\text{ and }x_{j}\in C_{i}\}$. For $k+1\leq i\leq m$, we have $\tilde{x}(i):=\min\{j\colon j\in A^{C}\text{ and }\neg x_{j}\in C_{i}\}$. The order in which trains traverse shared edges is not predefined; hence, w.l.o.g. assume that $\tilde{x}(i)\geq\tilde{x}(i+1)$ for every $11\leq i\leq k-$ 1 and $\tilde{x}(i)\leq\tilde{x}(i+1)$ for every $k+1\leq i\leq m-1$. Otherwise, we can rename the trains at this point. For $1\leq i\leq m$, we let $\operatorname{tr}_{i}$ stop on $(c_{in}^{(1)},x_{0})$ until $t=(i-1)\cdot T_{n}+1$. Afterwards, trains $i=1,\ldots,k$ travel at maximal speed all the way to $(x_{\tilde{x}(i)-1,\tilde{x}(i)}^{(r)},x_{\tilde{x}(i)}^{(1)})$ or $(x_{\tilde{x}(i)-1}^{(3)}),x_{\tilde{x}(i)-1,\tilde{x}(i)}^{(r)})$ if the previous edge is already occupied. Because a literal appears in only two clauses, it cannot happen that both edges are already occupied. Similarly, for trains traveling in the reverse direction ($i=k+1,\ldots,m$), travel at maximal speed all the way to $(x_{\tilde{x}(i),\tilde{x}(i)+1}^{(l)},x_{\tilde{x}(i)}^{(3)})$ or $(x_{\tilde{x}(i)+1}^{(1)}),x_{\tilde{x}(i),\tilde{x}(i)+1}^{(l)})$ if the previous edge is already occupied. Again, because a literal appears in only two clauses, it cannot happen that both edges are already occupied. All trains will arrive latest at time $t=i\cdot T_{n}$. At time $t=m\cdot T_{n}$, they can all traverse two edges into stop $S$ within at most four time steps. Finally, they can proceed in the same order to the respective exit vertices in, again, at most $m\cdot T_{n}$ time steps and wait on their last edge without affecting other trains. There are no conflicts because of the chosen order and feasibility of the monotone 3-SAT-(2,2$\leq$ 2$\leq$ 2≤ 2 , ≤ 2). In particular, this constitutes a feasible solution to the corresponding TS-VER instance. In Figure 14, the trains are depicted at $t_{0}=1$, $t_{1}=m\cdot T_{n}$, and $t_{2}=m\cdot T_{n}+5$. This setting corresponds to setting $x_{1}$ to false and $x_{4}$ to true.
@@
On the other hand, if we are given a feasible timing of the corresponding TS-VER instance, we check the positions at time $t=m\cdot T_{n}+5$. We set $x_{j}$ to be true if $(x_{j}^{(1)},x_{j}^{2})$ or $(x_{j}^{(2)},x_{j}^{3})$ is occupied and to false if $(x_{j}^{(3)},x_{j}^{2})$ or $(x_{j}^{(2)},x_{j}^{1})$ is occupied. All other variables are chosen arbitrarily. Note that a variable cannot be set to true and false because then trains would collide on the opposing tracks. Hence, this constitutes a feasible solution to the corresponding monotone 3-SAT-(2,2$\leq$ 2$\leq$ 2≤ 2 , ≤ 2) instance by design.
@@
Combining the above, we have argued that a monotone 3-SAT-(2,2$\leq$ 2$\leq$ 2≤ 2 , ≤ 2) instance is feasible if, and only if, the corresponding TS-VER instance is. Hence, we have polynomially reduced Monotone 3-SAT-(2,2$\leq$ 2$\leq$ 2≤ 2 , ≤ 2) to TS-VER, proving the latter’s NP-completeness. ∎
@@ III-B Generation of VSS Layouts
By leaving the TTDs fixed and only adding VSS sections, we can change the block structure at a low cost. Because of this, it might be possible to change the virtual subsections when a new train schedule is introduced. Hence, we can somewhat “overfit” to the current schedule without relying solely on schedule-independent KPIs (e.g., headway). This leads to the following design task of generating VSS layouts. {problem} If possible, we try to keep the number of additional virtual subsections small to keep it tractable for the control system during operation. In particular, we do not require a hard optimum because one section more or less will likely not pose any additional costs. On the other hand, it might be interesting to still have a minimal number of subsections to ensure a given schedule. In a later stage, one could then add more subsections to improve robustness; if the generated VSS layout from this design task is minimal, there might be more space to add subsections to improve robustness.
@@
Again, consider the setting from Example III.1 and Figure 13. If we separate $\tilde{e}_{10}$ into two virtual subsections by including a VSS border ( ), the train station $S_{0}$ has three blocks. In fact, we can now route the trains as shown in Figure 16, i.e., in the end, trains $\operatorname{tr}_{1}$ and $\operatorname{tr}_{2}$ can occupy the newly created subsections and train $\operatorname{tr}_{3}$ the other track. Hence, by adding one VSS border, we have made the previously infeasible schedule feasible. Note that we could also have added additional subsections, e.g., split track $\tilde{e}_{5}$ and have $\operatorname{tr}_{3}$ occupy two blocks in the end. However, as said, our objective is to add rather few VSS borders, and one additional VSS border suffices to make the schedule feasible.
@@
Consider an instance of TS-VER. Assume we are given an oracle to solve any instance of VSS-GEN. The given TS-VER instance constitutes valid input to VSS-GEN. Let $k^{\star}$ be the number of VSS addition operators used in an optimal solution of VSS-GEN on the respective input or let $k^{\star}=\infty$ if no feasible solution exists. Then, the solution to the given TS-VER instance is “yes” if, and only if, $k^{\star}=0$. In particular, we have polynomially reduced TS-VER to VSS-GEN. The claim follows directly from Theorem III.2. ∎
@@ III-C Schedule Optimization Using the Potential of VSS
In practice, minimizing VSS sections might not be of great interest. Instead, one might have several VSS sections that can be safely handled. In this case, one wants to place them strategically to optimize the operational outcome, e.g., to minimize travel times. {problem}
Another operational objective might be to maximize capacity. Assume, e.g., one plans a mixed railway network on which passenger trains should operate, adhering to certain scheduling constraints. On the other hand, there are additional requests for freight trains running through the network. In this setting, one might want to maximize the number of freight trains without negatively affecting the schedule of passenger trains by using the full potential of VSS sections.
@@
VSS-OPT and VSS-CAP are NP-hard even if all (potential) routes and exact stopping times in train stations are fixed, and the railway network is planar
@@
Consider an instance of TS-VER. Assume we are given an oracle to solve any instance of VSS-OPT. The given TS-VER instance constitutes valid input to VSS-OPT. Let $t^{\star}$ be the optimal travel time objective or $t^{\star}=\infty$ if no feasible solution exists. Observe that it is easy to deduce a theoretical upper bound on $t^{\star}$ in the feasible case, say $\overline{t}$, from the latest possible exit times $\overline{t}_{out}^{tr_{1}},\ldots,\overline{t}_{out}^{tr_{m}}$, e.g., $\overline{t}=\overline{t}_{out}^{tr_{1}}+\ldots+\overline{t}_{out}^{tr_{m}}$ if the objective is given by the sum of travel times. Then, the solution to the given TS-VER instance is “yes” if, and only if, $t^{\star}\neq\infty$, i.e., $t^{\star}\leq\overline{t}$.
The given TS-VER instance also constitutes valid input to VSS-CAP. Let $r^{\star}=|\mathcal{R^{\star}}|$ be the maximal number of optional trains $\tilde{\operatorname{tr}}_{1},\ldots,\tilde{\operatorname{tr}}_{m_{2}}$ that are routed or $r^{\star}=-\infty$ if the problem is infeasible even for $\mathcal{R}=\emptyset$. Then, the solution to the given TS-VER instance is “yes” if, and only if, $r^{\star}\neq\infty$, i.e., $r^{\star}\leq m_{2}$.
In particular, we have polynomially reduced TS-VER to VSS-OPT and VSS-CAP. The claim follows directly from Theorem III.2. ∎
@@ IV Conclusions
The results confirm that previously proposed heuristics and exact methods (presented, e.g., in [34, 35]), which only generate non-optimal results or have a severely limited scalability, respectively, do not suffer from insufficient methodology but complexity reasons. Overall, for the first time, a solid basis for developing (ideally scalable) design automation solutions incorporating the freedom allowed by virtual subsections within ETCS L2 HTD is provided. These methods might be a valuable addition to the already existing work on fixed-block control systems summarized in Section I. It will also be interesting to see how these methods tailored to hybrid train detection systems perform compared to previously designed methods that do not consider the specific schedule for generating signaling layouts.

### v5

@@ III-A Verification of Timetable Requests on ETCS L2 HTD Layouts
Assume that the given monotone 3-SAT-(2,2$\leq$ 2$\leq$ 2≤ 2 , ≤ 2) instance is feasible. From any particular solution, define $A:=\{j\colon x_{j}\text{ is true}\}$ and $A^{C}:=\{1,\ldots,m\}-A=\{j\colon x_{j}\text{ is false}\}$. For $1\leq i\leq k$, define $\tilde{x}(i):=\min\{j\colon j\in A\text{ and }x_{j}\in C_{i}\}$. For $k+1\leq i\leq m$, we have $\tilde{x}(i):=\min\{j\colon j\in A^{C}\text{ and }\neg x_{j}\in C_{i}\}$. The order in which trains traverse shared edges is not predefined; hence, w.l.o.g. assume that $\tilde{x}(i)\geq\tilde{x}(i+1)$ for every $11\leq i\leq k-$ 1 and $\tilde{x}(i)\leq\tilde{x}(i+1)$ for every $k+1\leq i\leq m-1$. Otherwise, we can rename the trains at this point. For $1\leq i\leq m$, we let $\operatorname{tr}_{i}$ stop on $(c_{in}^{(1)},x_{0})$ until $t=(i-1)\cdot T_{n}+1$. Afterwards, trains $i=1,\ldots,k$ travel at maximal speed all the way to $(x_{\tilde{x}(i)-1,\tilde{x}(i)}^{(r)},x_{\tilde{x}(i)}^{(1)})$ or $(x_{\tilde{x}(i)-1}^{(3)}),x_{\tilde{x}(i)-1,\tilde{x}(i)}^{(r)})$ if the previous edge is already occupied. Because a literal appears in only two clauses, it cannot happen that both edges are already occupied. Similarly, for trains traveling in the reverse direction ($i=k+1,\ldots,m$), travel at maximal speed all the way to $(x_{\tilde{x}(i),\tilde{x}(i)+1}^{(l)},x_{\tilde{x}(i)}^{(3)})$ or $(x_{\tilde{x}(i)+1}^{(1)}),x_{\tilde{x}(i),\tilde{x}(i)+1}^{(l)})$ if the previous edge is already occupied. Again, because a literal appears in only two clauses, it cannot happen that both edges are already occupied. All trains will arrive latest at time $t=i\cdot T_{n}$. At time $t=m\cdot T_{n}$, they can all traverse two edges into stop $S$ within at most four time steps. Finally, they can proceed in the same order to the respective exit vertices in, again, at most $m\cdot T_{n}$ time steps and wait on their last edge without affecting other trains. There are no conflicts because of the chosen order and feasibility of the monotone 3-SAT-(2,2$\leq$ 2$\leq$ 2≤ 2 , ≤ 2). In particular, this constitutes a feasible solution to the corresponding TR-VER instance. In Figure 14, the trains are depicted at $t_{0}=1$, $t_{1}=m\cdot T_{n}$, and $t_{2}=m\cdot T_{n}+5$. This setting corresponds to setting $x_{1}$ to false and $x_{4}$ to true.
@@
On the other hand, if we are given a feasible timing of the corresponding TR-VER instance, we check the positions at time $t=m\cdot T_{n}+5$. We set $x_{j}$ to be true if $(x_{j}^{(1)},x_{j}^{2})$ or $(x_{j}^{(2)},x_{j}^{3})$ is occupied and to false if $(x_{j}^{(3)},x_{j}^{2})$ or $(x_{j}^{(2)},x_{j}^{1})$ is occupied. All other variables are chosen arbitrarily. Note that a variable cannot be set to true and false because then trains would collide on the opposing tracks. Hence, this constitutes a feasible solution to the corresponding monotone 3-SAT-(2,2$\leq$ 2$\leq$ 2≤ 2 , ≤ 2) instance by design.
@@
Combining the above, we have argued that a monotone 3-SAT-(2,2$\leq$ 2$\leq$ 2≤ 2 , ≤ 2) instance is feasible if, and only if, the corresponding TR-VER instance is. Hence, we have polynomially reduced Monotone 3-SAT-(2,2$\leq$ 2$\leq$ 2≤ 2 , ≤ 2) to TR-VER, proving the latter’s NP-completeness. ∎
@@ III-B Generation of VSS Layouts
Some elements (e.g., turnouts) require special treatment and classical train separation. In our formulation, this can be achieved by marking some of the TTD sections as unbreakable. It is then forbidden to apply the VSS addition operator on an edge within any unbreakable TTD section. For example, any TTD section containing a turnout could be marked in such a way. By doing this, the control logic around switches would not change, while at the same time allowing optimization using VSS on all permitted tracks (only).
By leaving the TTDs fixed and only adding VSS sections, we can change the block structure at a low cost. Because of this, it might be possible to change the virtual subsections when a new train schedule should be implemented. Hence, we can somewhat “overfit” to the current timetable requests without relying solely on schedule-independent KPIs (e.g., headway). This leads to the following design task of generating VSS layouts. {problem} If possible, we try to keep the number of additional virtual subsections small to keep it tractable for the control system during operation. In particular, we do not require a hard optimum because one section more or less will likely not pose any additional costs. On the other hand, it might be interesting to still have a minimal number of subsections to make given timetable requests realizable. In a later stage, one could then add more subsections to improve robustness; if the generated VSS layout from this design task is minimal, there might be more space to add subsections to improve robustness.
@@
Again, consider the setting from Example III.1 and Figure 13. If we separate $\tilde{e}_{10}$ into two virtual subsections by including a VSS border ( ), the train station $S_{0}$ has three blocks. In fact, we can now route the trains as shown in Figure 16, i.e., in the end, trains $\operatorname{tr}_{1}$ and $\operatorname{tr}_{2}$ can occupy the newly created subsections and train $\operatorname{tr}_{3}$ the other track. Hence, by adding one VSS border, we have made the previously infeasible timetable requests realizable. Note that we could also have added additional subsections, e.g., split track $\tilde{e}_{5}$ and have $\operatorname{tr}_{3}$ occupy two blocks in the end. However, as said, our objective is to add rather few VSS borders, and one additional VSS border suffices to make the timetable requests realizable.
Figure 16: Generation of VSS layout $(t=300)$
@@
Consider an instance of TR-VER. Assume we are given an oracle to solve any instance of VSS-GEN. The given TR-VER instance constitutes valid input to VSS-GEN. Let $k^{\star}$ be the number of VSS addition operators used in an optimal solution of VSS-GEN on the respective input or let $k^{\star}=\infty$ if no feasible solution exists. Then, the solution to the given TR-VER instance is “yes” if, and only if, $k^{\star}=0$. In particular, we have polynomially reduced TR-VER to VSS-GEN. The claim follows directly from Theorem III.2. ∎
@@ III-C Schedule Optimization Using the Potential of VSS
In practice, minimizing VSS sections might not be of great interest. Instead, one might want to optimize the operational outcome, e.g., by minimizing travel times. Intuitively, one might think that this is an (computationally) easy problem, because the “best” layout is creating very small VSS (e.g., every meter or even less). In that case, trains would basically operate as if a moving block control system were implemented. However, according to industry experts, such an implementation would not be feasible because every virtual block has to be specified within the interlocking and the Radio Block Center (RBC). These components must fulfill the safety integrity level (SIL) 4 and are time-consuming and costly to develop and maintain. Because many such systems were developed and approved for classical control systems, the address space for block sections has limited capacity.
Thus, in practice, there is a significant upper limit on how many VSS can be defined, which should be considered within the planning process. Hence, the question arises on strategically placing these VSS to optimize the operational outcome. In particular, a proposed solution of placing a VSS every meter is likely not feasible.
{problem}
Another operational objective might be to maximize capacity. Assume, e.g., one plans a mixed railway network on which passenger trains should operate, adhering to certain scheduling constraints. On the other hand, there are additional requests for freight trains running through the network. In this setting, one might want to maximize the number of freight trains without negatively affecting the timetable requests of passenger trains by using the full potential of VSS sections.
@@
VSS-OPT and VSS-CAP are NP-hard even if all (potential) routes and exact stopping times in train stations are fixed, and the railway network is planar.
@@
Consider an instance of TR-VER. Assume we are given an oracle to solve any instance of VSS-OPT. The given TR-VER instance constitutes valid input to VSS-OPT. Let $t^{\star}$ be the optimal travel time objective or $t^{\star}=\infty$ if no feasible solution exists. Observe that it is easy to deduce a theoretical upper bound on $t^{\star}$ in the feasible case, say $\overline{t}$, from the latest possible exit times $\overline{t}_{out}^{tr_{1}},\ldots,\overline{t}_{out}^{tr_{m}}$, e.g., $\overline{t}=\overline{t}_{out}^{tr_{1}}+\ldots+\overline{t}_{out}^{tr_{m}}$ if the objective is given by the sum of travel times. Then, the solution to the given TR-VER instance is “yes” if, and only if, $t^{\star}\neq\infty$, i.e., $t^{\star}\leq\overline{t}$.
The given TR-VER instance also constitutes valid input to VSS-CAP. Let $r^{\star}=|\mathcal{R^{\star}}|$ be the maximal number of optional trains $\tilde{\operatorname{tr}}_{1},\ldots,\tilde{\operatorname{tr}}_{m_{2}}$ that are routed or $r^{\star}=-\infty$ if the problem is infeasible even for $\mathcal{R}=\emptyset$. Then, the solution to the given TR-VER instance is “yes” if, and only if, $r^{\star}\neq\infty$, i.e., $r^{\star}\leq m_{2}$.
In particular, we have polynomially reduced TR-VER to VSS-OPT and VSS-CAP. The claim follows directly from Theorem III.2. ∎
@@ IV Conclusions
The results confirm that previously proposed heuristics and exact methods (presented, e.g., in [35, 36]), which only generate non-optimal results or have a severely limited scalability, respectively, do not suffer from insufficient methodology but complexity reasons. Overall, for the first time, a solid basis for developing (ideally scalable) design automation solutions incorporating the freedom allowed by virtual subsections within ETCS L2 HTD is provided. These methods might be a valuable addition to the already existing work on fixed-block control systems summarized in Section I. It will also be interesting to see how these methods tailored to hybrid train detection systems perform compared to previously designed methods that do not consider the specific timetable requests for generating signaling layouts.